HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

HIGH-RESOLUTION FT-UV EMISSION SPECTROSCOPY OF THE 1-ν′′ BANDS OF B^{2}Σ^{+} − X^{2}Σ^{+} SYSTEM OF^{12}C^{16}O^{+}.

W. SZAJNA, R. KȨPA, \underline{R. HAKALLA} , A. PARA, I. PIOTROWSKA, S. RYZNER, A. STASIK,, Materials Spectroscopy Laboratory, Institute of Physics, University of Rzeszów, Pigonia 1 Street, 35-310 Rzeszów, Poland; M.I. MALICKA,

The Faculty of Mathematics and Applied Physics, Rzeszów University of Technology, Powstańców Warszawy 8 Street, 35-959 Rzeszów, Poland

The B^{2}Σ^{+} (ν = 1) level of^{12}C^{16}O^{+} was investigated using the high-resolution, emission spectra obtained via Fourier-transform spectroscopy of the strongest (1−2) and (1 − 3) bands of the First Negative (B^{2}Σ^{+} − X^{2}Σ^{+}) system^{1}^{,}^{2}.

The B-X, CO^{+} bands were recorded with an instrumental resolution of 0.04 cm^{−1} and the best signal-to-noise ratio ca. 50:1. For obtaining spectra of the ^{12}C^{16}O^{+} isotopologue an air-cooled hollow-cathode (HC) lamp, filled with^{12}C^{16}O_{2} and trace amount of Ne, was used. The spectrum was calibrated using measurements of the atomic lines of Fe ^{3} impurity present in the spectra. The absolute accuracy of the calibration is estimated to be better than 0.0050 cm^{−1}.

The present data were elaborated with the PGOPHER program^{4} and wide set of molecular constants for the B^{2}Σ^{+}, ν = 1 and X^{2}Σ^{+}, ν = 2, 3 and levels was obtained.This work results were compared with the recent ones presented by Ventura et al.^{5}, and the discussed by the authors band origin values were carefully checked and verified. Moreover, a very week irregularities with the positions of the lines connected with B^{2}Σ^{+}, ν = 1, N = 21 levels were discovered.

1doi:10.1016/j.jms.2003.10.007, W. Szajna, R. Kȩpa, M. Zachwieja, The first negative (B^{2}Σ^{+} − X^{2}Σ^{+}) system of CO^{+}: excitation of higher vibrational levels, J. Mol. Spectrosc., 223, 125-131 (2004)

2doi:10.1140/epjd/e2004-00060-0, W. Szajna, R. Kȩpa, M. Zachwieja, Reinvestigation of the B^{2}Σ^{+} − X^{2}Σ^{+} system in the CO^{+}, Eur. Phys J. D, 30, 49-55 (2004).

3doi:10.1364/JOSAB.8.002028, G. Nave, R.C.M. Learner, A.P. Thorne, C.J. Harris, Precision Fe I and Fe II wavelengths in the ultraviolet spectrum of the iron-neon hollow-cathode lamp, J. Opt. Soc.

Am. B, 8, 2028-2041, (1991).

4doi:10.1016/j.jqsrt.2016.04.010, C. Western, PGOPHER: A program for simulating rotational, vibrational and electronic spectra, J. Quant. Spectrosc. Radiat. Transfer, 186, 221-242 (2017).

5doi:10.1016/j.jqsrt.2020.107312, L.R. Ventura, R.S. da Silva, M.Y. Balleser, C.E. Fellows, Revisiting^{12}C^{16}O^{+} first negative system: An experimental and theoretical study, J. Quant. Spectrosc. Radiat.

Transfer, 256, 107312 (2020).

p-number: p075  Submitted on Mon, 12 Jun 2023 11:03:06 +0200